# Generative Modelling of Structurally Constrained Graphs

### v1

@@ 4.1 Synthetic Graphs
Regarding performance metrics, we follow the evaluation procedures from Martinkus et al. [53]. We assess how close the distributions of different graph statistics computed from the generated and test sets are. To accomplish that, we compute the Maximum Mean Discrepancy (MMD) [^2] To align with previous literature, we actually compute MMD2. for the node degrees (Deg.), clustering coefficients (Clus.), orbit count (Orbit), eigenvalues of the normalized graph Laplacian (Spec.), and statistics from a wavelet graph transform (Wavelet). To summarize this set of metrics, we compute the ratio over the same metrics obtained by the train set and average it (Ratio). We also compute the proportion of non-isomorphic graphs (Unique), the proportion of graphs non-isomorphic to any graph in the train set (Novel) and the proportion of valid generated graphs (Valid). Graphs are considered valid if they are planar and connected, trees, or lobster graphs, when the generative model is trained on the planar, tree, or lobster dataset, respectively. We merge these three metrics through the proportion of generated graphs that are simultaneously valid, unique and novel (V.U.N.).
@@
We present the results in Table 1. For the planar dataset, ConStruct achieves nearly optimal performance, clearly outperforming all other methods. It is actually the first method to achieve 100% V.U.N., indicating state-of-the-art performance. Moreover, in terms of average ratio, it clearly outperforms all other methods, with the average ratio approaching 1, suggesting high sample quality. Regarding the lobster dataset, ConStruct exhibits a similar trend, demonstrating superior performance compared to DiGress+. It leads to state-of-the-art results in both average ratio and V.U.N. metrics. The lower novelty and uniqueness values ($<$100%) are attributed to the dataset’s smaller size. In fact, we train both models on 64 examples (80% of the train set) while generating 100 graphs in each run. Conversely, for the tree dataset, ConStruct is outperformed by DiGress+ due to the marginally lower expressivity of the edge-absorbing noise model for this particular case (see Section H.1 for details). As a sanity check, we observe that for all three datasets, ConStruct ensures the constraining property for all generated graphs. However, the validity values are below 100% (except for planar) since connectedness of the graph is not guaranteed. In general, this property is not ensured by one-shot models and cannot be included as a constraining property since it is not edge-deletion invariant.
@@ 4.2 Digital Pathology Cell Graphs
In the next set of experiments, we explore digital pathology data. Due to the their natural representation of relational data, graphs are widely used to capture spatial biological dependencies from tissue images. We focus on cell graphs, whose nodes represent biological cells and edges serve as proxies for local cell-cell interactions. We build these structures from the genomic and clinical data available from the Molecular Taxonomy of Breast Cancer International Consortium (METABRIC) molecular dataset [12, 65, 14]. Each node is attributed with one of the nine possible phenotypes, which extensively characterizes a cell both anatomically and physiologically (more details in  Section F.2). Regarding edges, we followed the typical procedure for cell graphs in digital pathology [34, 35, 2, 80]: first we employ Delaunay triangulation on the cell positions to construct the graphs, followed by edge thresholding to discard long edges. Our focus lies on generating biologically meaningful Tertiary Lymphoid Structures (TLSs), further described in Section F.3. Thus, we extract non-overlapping 4-hop subgraphs centered at nodes whose class is “B” from the whole-slide graphs. In terms of dimensionality, we obtain graphs with $b=9$, corresponding to the 9 phenotypes detailed in Section F.2, and $c=1$. We explore two datasets: one comprising graphs with high TLS content and another consisting of graphs with low TLS content, based on domain-specific metrics (see below). We provide their statistic in Section F.4. We open-source both of them, representing to the best of our knowledge the first open-source digital pathology datasets specifically tailored for graph generation. For the sake of comparison, besides ConStruct and DiGress+, we implement a non deep learning baseline method proposed in [51] for this setting, which essentially captures 1-hop dependencies of cell graphs (see Section F.5). Additionally, we run BiGG [13], GraphGen [23], and SPECTRE [53]. These are the methods that, besides DiGress, attain non-zero V.U.N. for the planar dataset in Table 1, which we consider a proxy for performance in the digital pathology datasets due to the structural similarities between the datasets.
@@
The TLS embedding, $\kappa=[\kappa_{0},\ldots,\kappa_{5}]\in\mathbb{R}^{6}$, has been proposed to quantify the TLS content in a cell graph [67, 51]. See Section F.3 for more details. Based on this metric, we define a graph $G$ to contain low TLS content if $k_{1}(G)<0.05$ and with high TLS content $k_{2}(G)>0.05$ [51]. To evaluate the generative performance, we adopt the average ratio for structural graph statistics (Ratio) and V.U.N. metrics used in Section 4.1. Here we consider a planar and connected graph as a valid graph. Thus, we present explicitly the proportion of generated graphs that are connected (Conn.) and (Planar). Furthermore, for a biologically meaningful evaluation of the generated cell graphs, we use the domain metrics. We report the MMD between the distributions of the components of $\kappa$. We also consider the proportion of graphs that are planar, connected, and verify the low or high TLS content condition (TLS Valid), depending on the train set used.
@@ Results
ConStruct outperforms all the baselines for all the summary evaluation metrics (in light gray in Table 2) for cell graph generation in both datasets. Unlike the synthetic datasets, here the structural distribution is conditioned on the node types, which is inherently a more complex task. This complexity contributes to the poor performance of the several unconstrained models. Constraining the edge generation process allows to significantly alleviate this modelling complexity, highlighting the benefits of ConStruct in such scenarios. We emphasize the substantial improvement in the V.U.N. of the generated graphs, with values approaching 100% using our framework, which aligns with the main motivation behind the proposed method. Interestingly, it also promotes the generation of more connected graphs. Finally, the 1-hop baseline model, while capturing the node type dependencies to some extent, as illustrated by the MMD on the components of $\kappa$, fails completely to capture structure based dependencies.
Additionally, we carry out some experiments for molecular datasets in Appendix G: we explore the utilization of planarity for constrained molecular generation and showcase how ConStruct can be used for controlled generation. Finally, we explore likelihood-based variants of ConStruct as well as some ablations to the projector in Appendix H.
@@ 5 Conclusion
In this paper, we introduced ConStruct, a framework that allows to integrate domain knowledge via structural constraints into graph diffusion models. By constraining the diffusion process based on a diverse set of geometric properties, we enable the generation of realistic graphs in scenarios with limited data. To accomplish that, we leverage on an edge-absorbing noise model and a projector operator to ensure that both the forward and reverse processes maintain the validity of sampled graphs within the constrained domain. Despite its algorithmic simplicity, our approach overcomes the arbitrarily hard problem of projecting a given graph into a combinatorial subspace in an efficient and theoretically grounded manner. Through several experiments on benchmark datasets, we showcase the versatility of ConStruct across various structural constraints. For example, in digital pathology datasets, our method outperforms existing approaches, bringing the validity of the generated graphs close to 100%. The incorporation of constraining properties jointly dependent on node and edges, as well as further exploitation of the induced sparsity by the edge-absorbing noise model may offer promising extensions to ConStruct.
@@ Appendix A Graph Discrete Diffusion Model
A.1 Parameterization of the Reverse Process
$p_{\theta}(x_{i}^{t-1}|x_{i}=x,G^{t})=$
=$=\begin{cases}\frac{\mathbf{x}_{i}^{t}(\mathbf{Q}_{X}^{t})^{\prime}\odot\;\mathbf{x}_{i}\bar{\mathbf{Q}}^{t-1}_{X}}{\mathbf{x}_{i}^{t}\bar{\mathbf{Q}}^{t}_{X}\mathbf{x}_{i}}&\text{if }q(x_{i}^{t}|x_{i}=x)>0,\\ 0&\text{otherwise,}\end{cases}$ (6)
where denotes transposition.
To endow the graph diffusion model of the desired equivariance properties, a Graph Neural Network, $\text{GNN}_{\theta}$, is used as the denoising neural network, along with a permutation invariant loss (see Section A.2). In terms of architecture, we adopt the exact same denoising network architecture of DiGress [77], a Graph Transformer [19].
Importantly, the edge-absorbing noise model used in ConStruct increases graph sparsity throughout the forward trajectory. Consequently, beyond the distribution preserving guarantees, it also allows for the efficient computation of extra features on the noisy graphs that otherwise the $\text{GNN}_{\theta}$ would not be able to capture. Following Vignac et al. [77], these are fed as an additional input to the denoising network (see Algorithm 1 and Algorithm 2), further enhancing its expressivity beyond the well-known limited representational power of GNN architectures [82, 56]. More concretely, besides the spectral (eigenvalues and eigenvectors of the Laplacian) and structural (number of cycles) features from DiGress, we also consider some additional features. We add as graph features the degree distribution and the node and edge type distributions. While the former enhances the positional information within the graph, the latter helps in making more explicit to the model the prevalence of each class in the dataset. Additionally, we add auxiliary structural encodings to edges to boost edge label prediction. We compute the Adamic-Adar index to aggregate local neighborhood information and the shortest distance between nodes to encode node interactions. For computational limitations, we only consider information within a 10-hop radius for these computations. These additional features were previously proposed by Qin et al. [63].
A.2 Training Algorithm
The chosen denoising graph neural network $\text{GNN}_{\theta}$ is trained using the cross-entropy loss between its predicted probabilities for each node and edge types, $\hat{p}^{G}=(\hat{p}^{X},\hat{p}^{E})$ and the actual node and edge types of a clean graph, $\mathbf{G}=(\mathbf{X},\mathbf{E})$:
@@ A.2 Training Algorithm
This loss is permutation invariant. Thus, if we also consider an equivariant architecture (e.g. the one described in Section A.1), the diffusion model is endowed of the desired equivariance properties. This allows the model to dodge the node ordering sensitivity from which, for example, autoregressive models suffer.
Given these two components, all the necessary elements are in place for the training of the diffusion model, which is defined in Algorithm 1.
@@ A.3 Sampling Algorithm
In this section, we present Construct’s proposed sampling algorithm, shown in Algorithm 2. At each timestep, we use the projector operator, described in Algorithm 3, to keep the sampled graphs throughout the reverse process within the constrained domain.
Input: Number of graphs to sample $N$ and constraining property $P$
1 for $i=1$ to $N$ do
2       Sample $n$ from the training set distribution;
       Sample $G^{T}\sim q_{X}(n)\times q_{E}(n)$ ;
        // Sample from limit distribution
3       for $t=T$ to $1$ do
             $h\leftarrow f(G^{t},t)$ ;
              // Compute extra features
4             $\hat{p}^{X},\hat{p}^{E}\leftarrow\text{GNN}_{\theta}(G^{t},h)$;
             $\hat{G}^{t-1}\sim p_{\theta}(G^{t-1}|G^{t})$ ;
              // Sample from distribution in Equation 4
5             $G^{t-1}\leftarrow\operatorname{Projector}(P,\hat{G}^{t-1},G^{t})$;
6
7       end for
8      Store $G^{0}$;
9
10 end for
Input: Constraining property $P$, noisy graph $G^{t}=(X^{t},E^{t})$ and candidate graph $\hat{G}^{t-1}=(\hat{X}^{t-1},\hat{E}^{t-1})$
@@ Appendix B Theoretical Analysis
  • $P$ be the edge-deletion invariant Definition 3.1 constraining property of the projector;
@@
Now, since the reverse process of the diffusion model is an edge insertion process, we have $\hat{G}^{t-1}=G^{t}\cup E_{\text{candidate}}\supset G^{t}$. Also, we notice that the projector amounts to randomly remove edges from $\hat{G}^{t-1}$ until we find a graph within the constraint set (equivalently, it entails adding as many edges as possible to $G^{t}$ while ensuring that the graph remains within the constraint set). Thus, it suffices to prove that for $(G^{*},e^{*})$ we necessarily have an optimal edit path $e^{*}=(e^{*}_{1},...,e^{*}_{k})$ from $\hat{G}^{t-1}$ to $G^{*}$ exclusively composed of edge deletions. In this case, our projector can necessarily produce $G^{*}$.
@@
$|E_{\text{projector}}|=CC_{\text{candidate}}-1,$
which tightly matches the upper bound for the optimization problem seen above. Consequently, any $G_{\text{projector}}\in\mathcal{G}^{t-1}$ is necessarily solution of the optimization problem, i.e.:

### v2

@@ 4.1 Synthetic Graphs
Regarding performance metrics, we follow the evaluation procedures from Martinkus et al. [54]. We assess how close the distributions of different graph statistics computed from the generated and test sets are. To accomplish that, we compute the Maximum Mean Discrepancy (MMD) [^2] To align with previous literature, we actually compute MMD2. for the node degrees (Deg.), clustering coefficients (Clus.), orbit count (Orbit), eigenvalues of the normalized graph Laplacian (Spec.), and statistics from a wavelet graph transform (Wavelet). To summarize this set of metrics, we compute the ratios against the corresponding metrics from the training set and then average them (Ratio). We also compute the proportion of generated graphs that are non-isomorphic to each other (Unique), the proportion that are non-isomorphic to any graph in the training set (Novel), and the proportion of generated graphs that are valid (Valid). Graphs are considered valid if they are planar and connected, trees, or lobster graphs, when the generative model is trained on the planar, tree, or lobster dataset, respectively. We merge these three metrics through the proportion of generated graphs that are simultaneously valid, unique and novel (V.U.N.).
@@
We present the results in Table 1. For the planar dataset, ConStruct achieves nearly optimal performance, clearly outperforming all other methods. It is actually the first method to achieve 100% V.U.N., indicating state-of-the-art performance. Moreover, in terms of average ratio, it clearly outperforms all other methods, with the average ratio approaching 1, suggesting high sample quality. Regarding the lobster dataset, ConStruct exhibits a similar trend, demonstrating superior performance compared to DiGress+. It leads to state-of-the-art results in both average ratio and V.U.N. metrics. The lower novelty and uniqueness values ($<$100%) are attributed to the dataset’s smaller size. In fact, we train both models on 64 examples (80% of the train set) while generating 100 graphs in each run. Conversely, for the tree dataset, ConStruct is outperformed by DiGress+ due to the marginally lower expressivity of the edge-absorbing noise model for this particular case (see Section H.1 for details). As a sanity check, we observe that for all three datasets, ConStruct ensures the constraining property for all generated graphs. However, the validity values are below 100% (except for planar) since connectedness of the generated graph is not guaranteed. In general, this property is not ensured by one-shot models and cannot be included as a constraining property since it is not edge-deletion invariant.
@@ 4.2 Digital Pathology Cell Graphs
In the next set of experiments, we explore digital pathology data. Due to the their natural representation of relational data, graphs are widely used to capture spatial biological dependencies from tissue images. We focus on cell graphs, whose nodes represent biological cells and edges serve as proxies for local cell-cell interactions. We build these structures from the genomic and clinical data available from the Molecular Taxonomy of Breast Cancer International Consortium (METABRIC) molecular dataset [13, 66, 15]. Each node is attributed with one of the nine possible phenotypes, which extensively characterizes a cell both anatomically and physiologically (more details in  Section F.2). Regarding edges, we followed the typical procedure for cell graphs in digital pathology [36, 35, 2, 82]: first we employ Delaunay triangulation on the cell positions to construct the graphs, followed by edge thresholding to discard long edges. Our focus lies on generating biologically meaningful Tertiary Lymphoid Structures (TLSs), further described in Section F.3. Thus, we extract non-overlapping 4-hop subgraphs centered at nodes whose class is “B” from the whole-slide graphs. In terms of dimensionality, we obtain graphs with $b=9$, corresponding to the 9 phenotypes detailed in Section F.2, and $c=1$. We explore two datasets: one comprising graphs with high TLS content and another consisting of graphs with low TLS content, based on domain-specific metrics (see below). We provide their statistics in Section F.4. We open-source both of them, representing to the best of our knowledge the first open-source digital pathology datasets specifically tailored for graph generation. For the sake of comparison, besides ConStruct and DiGress+, we implement a non-deep learning baseline method proposed in [52] for this setting, which essentially captures 1-hop dependencies of cell graphs (see Section F.5). Additionally, we run BiGG [14], GraphGen [24], and SPECTRE [54]. These are the methods that, besides DiGress, can handle attributed graphs and attain non-zero V.U.N. for the planar dataset in Table 1, which we consider a proxy for performance in the digital pathology datasets due to the structural similarities (i.e., planarity) between the datasets.
@@
The TLS embedding, $\kappa=[\kappa_{0},\ldots,\kappa_{5}]\in\mathbb{R}^{6}$, has been proposed to quantify the TLS content in a cell graph [69, 52]. See Section F.3 for more details. Based on this metric, we define a graph $G$ to contain low TLS content if $k_{1}(G)<0.05$ and high TLS content if $k_{2}(G)>0.05$ [52]. To evaluate the generative performance, we adopt the average ratio for structural graph statistics (Ratio) and V.U.N. metrics used in Section 4.1. Here we consider a planar and connected graph as a valid graph. Thus, we explicitly present the proportion of generated graphs that are connected (Conn.) and (Planar). Furthermore, for a biologically meaningful evaluation of the generated cell graphs, we use the domain metrics. We report the MMD between the distributions of the components of $\kappa$. We also consider the proportion of graphs that are planar, connected, and verify the low or high TLS content condition (TLS Valid), depending on the train set used.
@@ Results
ConStruct outperforms all baselines across all summary evaluation metrics (shown in light gray in Table 2) for cell graph generation on both datasets. Unlike the synthetic datasets, here the structural distribution is conditioned on the node types, which is inherently a more complex task. This complexity contributes to the poor performance of the several unconstrained models. Constraining the edge generation process allows to significantly alleviate this modelling complexity, highlighting the benefits of ConStruct in such scenarios. We emphasize the substantial improvement in the V.U.N. of the generated graphs, with values approaching 100% using our framework, which aligns with the main motivation behind the proposed method. Interestingly, it also promotes the generation of more connected graphs. Finally, the 1-hop baseline model, while capturing the node type dependencies to some extent, as illustrated by the MMD on the components of $\kappa$, completely fails to capture structure-based dependencies.
Additionally, we carry out some experiments for molecular datasets in Appendix G: we explore the utilization of planarity for constrained molecular generation and showcase how ConStruct can be used for controlled generation. Finally, we explore likelihood-based variants of ConStruct, as well as some ablations to the projector in Appendix H.
5 Limitations and Future Directions
In our work, we cover edge-deletion invariant properties. However, ConStruct can be easily extended to also handle edge-insertion invariant properties (i.e., properties that hold upon edge insertion). This extension can be useful in domains where constraints such as having at least $n$ cycles in a graph are important. To achieve this, we can simply "invert" the proposed framework: design the transition matrices with the absorbing state in an existing edge state (instead of the no-edge state) and a projector that removes edges progressively (instead of inserting them) while conserving the desired property.
In the particular context of molecular generation, Appendix G illustrates that, while purely structural constraints can guide the generation of molecules with specific structural properties (e.g., acyclicity), for general properties shared by all molecules (e.g., planarity) they are too loose. In contrast, autoregressive models thrive in such setting due to the possibility of molecular node ordering (e.g., via canonical SMILES) and the efficient incorporation of joint node-edge constraints (e.g., valency). Therefore, although it consists of a fundamentally different setting than the one considered in this paper, incorporating joint node-edge constraints into ConStruct represents an exciting future direction.
Additionally, the induced sparsity created by the edge-absorbing noise model presents opportunities for further exploitation. By leveraging this sparsity, future extensions of ConStruct could enhance sampling efficiency and improve the underlying diffusion model’s scalability for generating larger graphs.
@@ 6 Conclusion
In this paper, we introduced ConStruct, a framework that allows to integrate domain knowledge via structural constraints into graph diffusion models. By constraining the diffusion process based on a diverse set of geometric properties, we enable the generation of realistic graphs in scenarios with limited data. To accomplish that, we leverage an edge-absorbing noise model and a projector operator to ensure that both the forward and reverse processes preserve the sampled graphs within the constrained domain and, thus, maintain their validity. Despite its algorithmic simplicity, our approach overcomes the arbitrarily hard problem of projecting a given graph into a combinatorial subspace in an efficient and theoretically grounded manner. Through several experiments on benchmark datasets, we showcase the versatility of ConStruct across various structural constraints. For example, in digital pathology datasets, our method outperforms existing approaches, bringing the validity of the generated graphs close to 100%. Overall, ConStruct opens new avenues for integrating domain-specific knowledge into graph generative models, thereby paving the way for their application in real-world scenarios.
Acknowledgements
We thank Cédric Vincent-Cuaz, Nikolaos Dimitriadis, Vaishnavi Subramanian, Yiming Qin, Sevda Öğüt, and Laura Toni for helpful discussions and feedback. We also thank Andreas Bergmeister and Yunhui Jang for helping set up the code to reproduce their experiments.
@@ Appendix A Graph Discrete Diffusion Model
A.1 Training Algorithm
The denoising neural network is trained using the cross-entropy loss between its predicted probabilities for each node and edge types, $\hat{p}^{G}=(\hat{p}^{X},\hat{p}^{E})$ and the actual node and edge types of a clean graph, $\mathbf{G}=(\mathbf{X},\mathbf{E})$:
@@ A.1 Training Algorithm
As shown by Vignac et al. [79], the loss in Equation 4 is node permutation invariant. Thus, if we also consider an equivariant architecture, the diffusion model is endowed of the desired equivariance properties, allowing the model to dodge the node ordering sensitivity from which, for example, autoregressive models suffer. For this reason, we adopt a Graph Neural Network, $\text{GNN}_{\theta}$, as the denoising neural network of our diffusion model. In particular, we employ the exact same denoising network architecture of DiGress [79], a Graph Transformer [20].
Importantly, the edge-absorbing noise model used in ConStruct increases graph sparsity throughout the forward trajectory. Consequently, beyond the distribution preserving guarantees, it also allows for the efficient computation of extra features on the noisy graphs that otherwise the $\text{GNN}_{\theta}$ would not be able to capture. Following Vignac et al. [79], these are fed as a supplementary input to the denoising network (see Algorithm 1 and Algorithm 3), further enhancing its expressivity beyond the well-known limited representational power of GNN architectures [84, 57]. More concretely, besides the spectral (eigenvalues and eigenvectors of the Laplacian) and structural (number of cycles) features from DiGress, we also consider some additional features. We add as graph features the degree distribution and the node and edge type distributions. While the former enhances the positional information within the graph, the latter helps in making more explicit to the model the prevalence of each class in the dataset. Additionally, we add auxiliary structural encodings to edges to boost edge label prediction. We compute the Adamic-Adar index to aggregate local neighborhood information and the shortest distance between nodes to encode node interactions. Due to computational limitations, we only consider information within a 10-hop radius for these computations. These additional features were previously proposed by Qin et al. [64].
Provided such loss function and denoising neural network architecture, all the necessary elements are in place for the training of the diffusion model, which is defined in Algorithm 1.
@@ 5       $\hat{p}^{X},\hat{p}^{E}\leftarrow\operatorname{GNN}_{\theta}(G^{t},h)$;
A.2 Parameterization of the Reverse Process
$p_{\theta}(x_{i}^{t-1}|x_{i}=x,G^{t})=\begin{cases}\frac{\mathbf{x}_{i}^{t}(\mathbf{Q}_{X}^{t})^{\prime}\odot\;\mathbf{x}_{i}\bar{\mathbf{Q}}^{t-1}_{X}}{\mathbf{x}_{i}^{t}\bar{\mathbf{Q}}^{t}_{X}\mathbf{x}_{i}}&\text{if }q(x_{i}^{t}|x_{i}=x)>0,\\ 0&\text{otherwise,}\end{cases}$ (7)
where denotes transposition.
@@ A.3 Sampling Algorithm
In this section, we first introduce the algorithm describing the proposed projector operator in Algorithm 2. This projector is employed at each time step to keep the sampled graphs throughout the reverse process within the constrained domain. The full sampling algorithm is shown in Algorithm 3.
Input: Constraining property $P$, noisy graph $G^{t}=(X^{t},E^{t})$, and candidate graph $\hat{G}^{t-1}=(\hat{X}^{t-1},\hat{E}^{t-1})$
@@ 7      $E^{\prime}\leftarrow E^{\prime}\setminus\{e^{\prime}\}$;
Input: Number of graphs to sample $N$ and constraining property $P$
1 for $i=1$ to $N$ do
       Sample $n$ from the training set distribution ;
        // Sample number of nodes
       Sample $G^{T}\sim q_{X}(n)\times q_{E}(n)$ ;
        // Sample from limit distribution
2       for $t=T$ to $1$ do
             $h\leftarrow f(G^{t},t)$ ;
              // Compute extra features
3             $\hat{p}^{X},\hat{p}^{E}\leftarrow\text{GNN}_{\theta}(G^{t},h)$;
             $\hat{G}^{t-1}\sim p_{\theta}(G^{t-1}|G^{t})$ ;
              // Sample from distribution in Equation 5
4             $G^{t-1}\leftarrow\operatorname{Projector}(P,\hat{G}^{t-1},G^{t})$;
5
6       end for
7      Store $G^{0}$;
8
9 end for
@@ Appendix B Theoretical Analysis
  • $P$ be the edge-deletion invariant (Definition 3.1) constraining property of the projector;
@@
Now, since the reverse process of the diffusion model is an edge insertion process, we have $\hat{G}^{t-1}=G^{t}\cup E_{\text{candidate}}\supset G^{t}$. Also, we notice that the projector amounts to randomly remove the edges that are not in $G^{t}$ from $\hat{G}^{t-1}$ until we find a graph within the constraint set (equivalently, it entails adding as many edges as possible to $G^{t}$ while ensuring that the graph remains within the constraint set). Thus, it suffices to prove that for $(G^{*},e^{*})$ we necessarily have an optimal edit path $e^{*}=(e^{*}_{1},...,e^{*}_{k})$ from $\hat{G}^{t-1}$ to $G^{*}$ exclusively composed of edge deletions. In this case, our projector can necessarily produce $G^{*}$.
@@
$|E_{\text{projector}}|\geq|\text{CC}_{\text{candidate}}|-1,$
which tightly matches the upper bound for the optimization problem seen above. Consequently, $|E_{\text{projector}}|=|\text{CC}_{\text{candidate}}|-1$ and any $G_{\text{projector}}\in\mathcal{G}^{t-1}$ is necessarily solution of the optimization problem, i.e.:
@@
Overall, Theorem 1, Theorem 2, and the counter-examples in Figure 3 show that the problem that the projector is addressing is not so trivial such that it can always output the optimal graph in the $\operatorname{GED}$ sense for all the edge-deletion invariant properties. Nevertheless, our projector is still guaranteed to produce graphs that meet the specified structural constraints.